OpenStudy (anonymous):

Medal A pharmacist receives an IV order from a doctor for 50 mg of cefaclor. The label on the bottle of cefaclor has a concentration of 25 mg of drug per 250 mL of liquid. How many mL of the liquid should the pharmacist prepare for the IV? ? mL

OpenStudy (anonymous):

there is \[\frac{ \text {25 mg of drug} }{\text{ 250 ml of liquid } }\] So using the same proportions How man ml of liquid would we need for 50 mg of the drug Notice that 25*2=50

OpenStudy (anonymous):

so the answer is 5?

OpenStudy (anonymous):

How did you get 5?

OpenStudy (anonymous):

25 / 250= 0.1 * 50 = 5

OpenStudy (anonymous):

25/250=.1 Means that 10% of the mixture is drugs so that means 50/x=.1 Makes no difference how many mg of drug we have we will always have the same propotion. 10% of the mixture will be drugs and 90% will be liquid Solve for x 50/x=.1

OpenStudy (anonymous):

so then answer is 1?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\frac{50}{x}=.1\] Solve for x